What is the mass (in grams) of the upper leg bone (femur) if it has a volume of 115 cm3 ?

Bone has a density of 1.8 g/cm3.


1
Expert's answer
2021-05-17T04:45:39-0400

"Density = \frac{m}{V} \\

1.8 = \frac{m}{115} \\

m = 1.8 \times 115 = 207 \;g"

Answer: 207 g